A tank initially contains 50 gal of fresh water. at t=0 brine solution containing 2 lb of salt per gallon is poured into the tank at the rate of 12 gal/minute. while the well-stirred mixtures leaves the tank at the rate of 8 gal/minutes. what is the amount of salt at the end of the 5 minutes?


1
Expert's answer
2021-12-03T14:50:02-0500
"\frac{ds}{dt}=24-\frac{8s}{50+4t}\\t=2.47\ min"